# Uncertainty Visualization via Low-Dimensional Posterior Projections

### v1

@@ 3.3 Architecture
Figure 6: MNIST inpainting and denoising. (a),(b) Application of PPDE to image inpainting from only the 8 bottom rows. The areas within the contours correspond to $50\%,80\%,90\%,98\%$ of the total probability mass. Note how the PPD reveals posterior multi-modality (e.g., it shows the digit is either a “7” or a “9”). Similarly, it also reveals intra-digit variations such as the circular part of the “6”. (c) Application of PPDE to image denoising with an extreme noise level of $\sigma_{\varepsilon}=1$. PPDE reveals a similar posterior multi-modality (e.g., the digit is either an “8” or a “5”).
4 Experiments
@@
We now illustrate our method on several tasks and datasets. In all our experiments with visual data, the subspace $\mathcal{A}(\bm{y})$ corresponds to the top two PCs of the posterior, obtained from NPPC [49]. Full details regarding the architectures and the per-task settings are in the Appendix.
4.1 Toy example in 2D
@@
Figure 5 illustrates the application of PPDE to a 2D toy example with 1D projections. Here, $\mathbf{x}$ is sampled from a mixture of six Gaussians (arranged as a face), and $\mathbf{y}$ is a noisy version of $\mathbf{x}$. The prior distribution $p_{\mathbf{x}}(\bm{x})$ and an exemplar sample from the joint distribution $(\bm{x},\bm{y})\sim p_{\mathbf{x},\mathbf{y}}$ are presented in Fig. 4(a). For this simple case, the posterior distribution can be calculated analytically (see Appendix for the derivation) and is also a mixture of Gaussians (Fig. 4(b)). Note that for this illustration the posterior itself is two-dimensional, and therefore can be plotted and visualized in full. However, as explained earlier, in the general high-dimensional case, this is precisely the problem our method aims to solve by projection. Here, this example serves as a sanity check enabling us to benchmark our results against a known ground truth.
@@
To demonstrate our method, we selected an arbitrary 1D subspace on which we project and plot the posterior. The origin point and a direction defining this subspace are shown in Fig. 4(b). As can be seen in Fig. 4(c), the PPD on the selected 1D subspace (green line) can provide us with insights regarding the behavior of the full posterior, such as the fact that it is composed of three modes with different weights. In addition, Fig. 4(c) also shows the probability density of the high dimensional (2D) posterior along the 1D line, which we refer to as the sliced posterior (gray line). Comparing both plots, it is easy to see why the sliced posterior is less informative as not all posterior modes appear in the slice projection.
4.2 Handwritten Digits
@@
Figures 6 demonstrate PPDE on inpainting and denoising of handwritten digits from the MNIST dataset. In the inpainting task, we used a mask that covers the top $70\%$ of the image, and in denoising we added noise of standard deviation $\sigma_{\varepsilon}=1$. As can be seen, for both tasks, PPDE reveals the inherent bimodality of the posterior with two different possible digits being likely given the measurement $\bm{y}$. In addition, note that when the digit identity is easier to infer from $\bm{y}$, PPDE reveals uncertainty corresponding to intricate intra-digit variations. More examples are available in the Appendix.
(a) Inpainting.
(b) Super-resolution 8$\times$.
Figure 7: CelebA inpainting and 8$\times$ super-resolution. (a) Application of PPDE to eye inpainting on the CelebA-HQ dataset. Regions of higher likelihood correlate with more natural eyes reflecting a positive correlation between $p_{\mathbf{v}|\mathbf{y}}{(\bm{v}|\bm{y})}$ and perceived plausibility. (b) Similarly, in super-resolution sharper images (e.g., B) are more likely than blurry ones (e.g., D).
Figure 8: Colorization. (a) Application of PPDE to face colorization on the CelebA-HQ dataset. The PPD is comprised of multiple connected modes along the first PC revealing background color ambiguity. Additionally, implausible skin colors (such as D) are significantly less likely along the second PC. (b)-(d) Application of PPDE to natural image colorization on the ImageNet dataset. Here, PPDE reveals typical grass and dog colors, where again implausible colorizations are significantly less likely (e.g., D).
4.3 Faces
To test PPDE on faces, we trained on the CelebA-HQ dataset using the original split inherited from celebA [35], resulting in 24183 images for training, 2993 images for validation, and 2824 images for testing. Figure 7 presents results on $256\times 256$ face images. Here, we demonstrate PPDE on the task of image inpainting on CelebA-HQ, as well as on 8×8$\times$8 × super-resolution with a bicubic downsampling filter on a center cropped version of the full celebA adopted from SRFlow [38]. As can be seen, likely solutions in the PPD are more semantically coherent, corresponding to more meaningful variations of $\bm{x}_{0}(\bm{y})$. Fig. 8 also shows the application of PPDE to face image colorization. Here, the PPD reveals the unique structure of the posterior, manipulating the background color and reflecting dataset bias. Note the plausibility of the most likely projected image (i.e., the maximum a-posteriori estimate), which can be easily obtained with our method. More results and examples are provided in the Appendix.
4.4 Natural images
Finally, we tested our approach on the large-scale ImageNet 1K dataset [57] for the task of image colorization. Fig. 8 shows that the resulting PPD reveals typical colors of grass and dogs. The resulting distribution is not necessarily multimodal, however, far from being an axis-aligned Gaussian, showing the practical benefit of the PPD. This validates the generality of PPDE on a large-scale dataset with rich posteriors. More examples are available in the Appendix.
@@ 4.4 Natural images
Figure 9: PPD comparison on MNIST. On the left we plot the input measurement $\bm{y}$, and the input-adaptive subspace $\mathcal{A}(\bm{y})=\{\bm{x}_{0}(\bm{y}),\bm{w}_{1}(\bm{y}),\bm{w}_{2}(\bm{y})\}$. In the middle, we plot the estimated PPDs obtained from a Gaussian approximation using NPPC, from the KDE approach, and from our PPDE. The projected ground truth posterior sample $\bm{x}$ is marked by a blue cross, and its likelihood under the estimated PPD is shown on the bottom left of the corresponding plot. On the right, we show a few posterior samples out of the 100 used in the KDE approach. These were obtained using an EBM trained on MNIST.
@@
Figure 10: PPD comparisons on CelebA-HQ. The layout is the same as Fig. 9. The posterior samples on the right were obtained using DDNM.
4.5 Quantitative comparisons
Comparison to NPPC
@@ 4.5 Quantitative comparisons
NPPC outputs the posterior mean $\bm{x}_{0}(\bm{y})$, alongside the top $K$ posterior PCs $\bm{W}(\bm{y})$ and variances $\{\sigma_{k}^{2}(\bm{y})\}_{k=1}^{K}$ for a given input image $\bm{y}$. While it was originally proposed as a distribution-free method, the predicted variances can be used to construct a 2D Gaussian approximation of the projected posterior covariance. Let $\bm{\Sigma}(\bm{y})$ denote the diagonal matrix that has $[\sigma_{1}^{2}(\bm{y}),\sigma_{2}^{2}(\bm{y})]^{\top}$ along its main diagonal, the projected posterior distribution can be approximated as
$p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})\approx\mathcal{N}(\bm{v};0,\bm{\Sigma}(\bm{y})).$ (9)
@@
We compare our results to this baseline in Table 1, showing a significant improvement in sample NLL in favor of PPDE.
@@
Table 1: Negative log likelihood comparison. NPPC and our method were computed on the entire test set. The KDE approach was computed for 100 test images due to computational constraints.
@@
Comparsion to sampling + KDE
The aim of our density estimator is to estimate the PPD in one or two dimensions. Given a method that can sample from the posterior (e.g., DDNM [72]), a possible baseline approach to achieve this goal is as follows: (i) Generate $N$ samples from the posterior $\{\bm{x}_{i}\}$; (ii) Compute PCA using $\{\bm{x}_{i}\}$ and project the samples onto the space spanned by the top two PCs; (iii) Estimate the PPD in 2D from the projected samples $\{\bm{v}_{i}\}$ using standard techniques such as kernel density estimation (KDE). Comparisons against this baseline are reported in Table 1. The posterior samplers used for this experiment are a conditional EBM for MNIST denoising/inpainting, MAT [33] for face inpainting, and DDNM for face colorization. For each task/dataset we randomly chose 100 test images and sampled 100 posterior samples per image. Compared to this (impractical) baseline, our technique is orders of magnitude faster as this approach requires costly repeated posterior sampling to estimate the projected density with enough accuracy. Figures 9 and  10 present visual examples of the approximated PPD with NPPC, KDE, and PPDE. PPDE exhibits superior NLL on average while being extremely faster than KDE, and only slightly slower than NPPC.
5 Discussion and conclusion
@@
Appendix A Interactive examples, code, and pretrained models
A website containing interactive examples and a link to the code and the trained models can be found at https://anonymous-monkey.github.io/PPDE/
Appendix B Toy model
B.1 Problem setting
@@
$p_{\mathbf{x}}(\bm{x})=\sum_{\ell=1}^{L}\pi_{\ell}\cdot\varphi(\bm{x};\bm{\mu}_{\ell},\bm{\Sigma}_{\ell}),$ (10)
@@
$\begin{array}[]{lll}\bm{\mu}_{1}=(-1,1)^{T}&\Sigma_{1}=\begin{pmatrix}0.05&-0.01\\ -0.01&0.025\end{pmatrix}&\pi_{1}=0.15\\ \bm{\mu}_{2}=(1,1)^{T}&\Sigma_{2}=\begin{pmatrix}0.05&0.01\\ 0.01&0.025\end{pmatrix}&\pi_{2}=0.15\\ \bm{\mu}_{3}=(0,0)^{T}&\Sigma_{3}=\begin{pmatrix}0.02&0\\ 0&0.03\end{pmatrix}&\pi_{3}=0.15\\ \bm{\mu}_{4}=(-0.7,-1)^{T}&\Sigma_{4}=\begin{pmatrix}0.15&-0.04\\ -0.04&0.04\end{pmatrix}&\pi_{4}=0.15\\ \bm{\mu}_{5}=(0,-1.2)^{T}&\Sigma_{5}=\begin{pmatrix}0.15&0\\ 0&0.025\end{pmatrix}&\pi_{5}=0.15\\ \bm{\mu}_{6}=(0.7,-1)^{T}&\Sigma_{6}=\begin{pmatrix}0.15&0.04\\ 0.04&0.04\end{pmatrix}&\pi_{6}=0.15.\\ \end{array}$ (11)
@@
B.2 Analytical posterior distribution
Assuming the setting in Sec. B.1, here we derive the analytical posterior distribution. Specifically, we show that the posterior is also a GMM, and that it has the form
$p_{\mathbf{x}|\mathbf{y}}(\bm{x}|\bm{y})=\sum_{\ell=1}^{L}\tilde{\pi}_{\ell}(\bm{y})\cdot\varphi(\bm{x};\tilde{\bm{\mu}}_{\ell}(\bm{y}),\tilde{\bm{\Sigma}}_{\ell}(\bm{y})),$ (12)
@@ B.2 Analytical posterior distribution
$\tilde{\pi}_{\ell}(\bm{y})$ $=\frac{\pi_{\ell}\varphi(\bm{y};\bm{\mu}_{\ell},\bm{\Sigma}_{\ell}+\sigma_{n}^{2}\bm{I})}{\sum_{\ell^{\prime}=1}^{L}\pi_{\ell^{\prime}}\varphi(\bm{y};\bm{\mu}_{\ell^{\prime}},\bm{\Sigma}_{\ell^{\prime}}+\sigma_{n}^{2}\bm{I})}.$ (13)
@@
$p_{\mathbf{x}}(\bm{x})=\sum_{\ell=1}^{L}p_{\mathbf{x}|\textnormal{c}}(\bm{x}|\ell)p_{\textnormal{c}}(\ell).$ (14)
@@
$p_{\mathbf{x}|\mathbf{y}}(\bm{x}|\bm{y})=\sum_{\ell=1}^{L}p_{\mathbf{x}|\mathbf{y},\textnormal{c}}(\bm{x}|\bm{y},\ell)p_{\textnormal{c}|\mathbf{y}}(\ell|\bm{y}).$ (15)
The term $p_{\textnormal{c}|\mathbf{y}}(\ell|\bm{y})$ in Eq. (15) can be written using Bayes’ rule as
@@
$=\frac{\pi_{\ell}\varphi(\bm{y};\bm{\mu}_{\ell},\bm{\Sigma}_{\ell}+\sigma_{n}^{2}\bm{I})}{\sum_{\ell^{\prime}=1}^{L}\pi_{\ell^{\prime}}\varphi(\bm{y};\bm{\mu}_{\ell^{\prime}},\bm{\Sigma}_{\ell^{\prime}}+\sigma_{n}^{2}\bm{I})}.$ (16)
Moreover, to explicitly express $p_{\mathbf{x}|\mathbf{y},\textnormal{c}}(\bm{x}|\bm{y},\ell)$ in Eq. (15), we note that $\mathbf{x}$ and $\mathbf{y}$ are jointly Gaussian given c, because under a particular choice $\textnormal{c}=\ell$, the random vector y is the sum of two independent Gaussian vectors, $\mathbf{x}$ and $\mathbf{n}$. Hence, conditioned on the event $\textnormal{c}=\ell$, the joint distribution of $\mathbf{x}$ and $\mathbf{y}$ is given by
$p_{\mathbf{x},\mathbf{y}|\textnormal{c}}(\bm{x},\bm{y}|\ell)$ $=\mathcal{N}\left(\begin{pmatrix}\bm{\mu}_{\ell}\\ \bm{\mu}_{\ell}\end{pmatrix},\begin{pmatrix}\bm{\Sigma}_{\mathbf{x}\mathbf{x}|\textnormal{c}}(\ell)&\bm{\Sigma}_{\mathbf{x}\mathbf{y}|\textnormal{c}}(\ell)\\ \bm{\Sigma}_{\mathbf{x}\mathbf{y}|\textnormal{c}}(\ell)&\bm{\Sigma}_{\mathbf{y}\mathbf{y}|\textnormal{c}}(\ell)\end{pmatrix}\right)$ (21)
$=\mathcal{N}\left(\begin{pmatrix}\bm{\mu}_{\ell}\\ \bm{\mu}_{\ell}\end{pmatrix},\begin{pmatrix}\bm{\Sigma}_{\ell}&\bm{\Sigma}_{\ell}\\ \bm{\Sigma}_{\ell}&\bm{\Sigma}_{\ell}+\sigma_{n}^{2}\bm{I}\end{pmatrix}\right).$ (26)
From the joint distribution in Eq. (26), the conditional distribution of $\mathbf{x}$ given $\mathbf{y}$ and c is derived as
$p_{\mathbf{x}|\mathbf{y},\textnormal{c}}(\bm{x}|\bm{y},\ell)=\varphi(\bm{x};\tilde{\bm{\mu}}_{\ell}(\bm{y}),\tilde{\bm{\Sigma}}_{\ell}(\bm{y})),$ (27)
with $\tilde{\bm{\mu}}_{\ell}(\bm{y})$ and $\tilde{\bm{\Sigma}}_{\ell}(\bm{y})$ defined as in Eq. (B.2).
B.3 Projected posterior distribution
Given that the posterior distribution is a GMM as in Eq. (12), it is straightforward to derive the projected posterior distribution (PPD) over any arbitrary 1D subspace defined by a center $\bm{x}_{0}$ and a normalized direction $\bm{w}$. This is done by projecting each of the Gaussians individually, i.e., the PPD for $v=\bm{w}^{T}(\bm{x}-\bm{x}_{0})$ is given by
$p_{\textnormal{v}|\mathbf{y}}(v|\bm{y})=\sum_{\ell=1}^{L}\tilde{\pi}_{\ell}(\bm{y})\cdot\varphi(v;\bm{w}^{T}(\tilde{\bm{\mu}}_{\ell}(\bm{y})-\bm{x}_{0}),\bm{w}^{T}\tilde{\bm{\Sigma}}_{\ell}(\bm{y})\bm{w}).$ (28)

### v2

@@ D Experiments
Figure S6: MNIST inpainting and denoising. (a),(b) Application of PPDE to image inpainting from only the 8 bottom pixel rows of the image. The areas within the contours correspond to $50\%,80\%,90\%,98\%$ of the total probability mass. Note how the PPD reveals posterior multi-modality (e.g., it shows the digit is either a “7” or a “9”). Similarly, it also reveals intra-digit variations such as the circular part of the digit “6”. (c) Application of PPDE to image denoising with an extreme noise level of $\sigma_{\varepsilon}=1$. PPDE reveals a similar posterior multi-modality (e.g., the digit is either an “8” or a “5”).
(a) Colorization
(b) Inpainting
Figure S7: Restoration of face images from CelebA-HQ. The PPD reveals a variety of plausible background and skin colors in colorization (a), and a range of plausible eye-shapes in inpainting (b). In addition, the PPD shows the likelihoods of the different restorations, distinguishing between plausible and less plausible solutions (such as points A and D, respectively, in (a)).
@@
We now illustrate our method on several tasks and datasets. In all our experiments with visual data, the subspace $\mathcal{A}(\bm{y})$ is spanned by the top two PCs of the posterior, obtained from NPPC [48]. Full details about the architectures, the per-task settings and additional results are in the appendix.
D.1 Toy example in 2D
@@
Figure S5 illustrates the application of PPDE to a 2D toy example with 1D projections. Here, $\mathbf{x}$ is sampled from a mixture of six Gaussians (arranged as a face), and $\mathbf{y}$ is a noisy version of $\mathbf{x}$. The prior distribution $p_{\mathbf{x}}(\bm{x})$ and an exemplar sample from the joint distribution $(\bm{x},\bm{y})\sim p_{\mathbf{x},\mathbf{y}}$ are presented in Fig. 5(a). For this simple case, the posterior distribution can be calculated analytically (see appendix for the derivation) and is also a mixture of Gaussians (Fig. 5(b)). Note that for this illustration the posterior itself is two-dimensional, and therefore can be plotted and visualized in full. However, as explained earlier, this is not the case for high-dimensional distributions and is precisely the problem our method aims to solve by projection. Here, this example serves as a sanity check enabling us to benchmark our results against a known ground truth.
@@
To demonstrate our method, we selected an arbitrary 1D subspace on which we project and plot the posterior. The origin point and a direction defining this subspace are shown in Fig. 5(b). As can be seen in Fig. 5(c), the PPD on the selected 1D subspace (green line) can provide us with insights regarding the behavior of the full posterior, such as the fact that it is composed of three modes with different weights. In addition, Fig. 5(c) also shows the probability density of the high dimensional (2D) posterior along the 1D line, which we refer to as the sliced posterior (purple line). Comparing both plots, it is easy to see why the sliced posterior is less informative as not all posterior modes appear in the slice projection.
D.2 Common restoration problems
Figure S8: Application of PPDE to natural image colorization on ImageNet, revealing the distribution of different solutions.
Figure S9: Biological image-to-image transfer. The PPD outlines the probability of possible solutions for the task of reconstructing fluorescent microscopy images using one type of dye from another. As seen in these examples, the PPD can assist in visualizing the fact that a given input might produce plausible solutions with varying amounts of cells (highlighted by yellow arrows).
Handwritten Digits.
@@ D.2 Common restoration problems
Figure S6 demonstrate PPDE on inpainting and denoising of handwritten digits from the MNIST dataset. In the inpainting task, we used a mask that covers the top $70\%$ of the image, and in denoising we added noise of standard deviation $\sigma_{\varepsilon}=1$. As can be seen, for both tasks, PPDE reveals the inherent bimodality of the posterior with two different possible digits being likely given the measurement $\bm{y}$. In addition, note that when the digit identity is easier to infer from $\bm{y}$, PPDE reveals uncertainty corresponding to intricate intra-digit variations.
Faces.
Figure S7 presents results for colorization and inpainting of face images from the CelebA-HQ $256\times 256$ dataset. In the appendix, we further provide results for 8×8$\times$8 × super-resolution with a bicubic downsampling filter and a crop area taken from SRFlow [37]. As can be seen, PPDE reveals the unique structure of the projected posterior distribution, which outlines the region of plausible solutions in the projected space.
Natural images.
Figure S8 shows colorization results for natural images from the ImageNet 1K dataset [56]. The distributions in this case are often not multi-modal; however, they are also far from being axis-aligned Gaussians. This demonstrate the practical benefit of PPD visualizations. More examples are available in the Appendix.
@@
Figure S10: PPD comparison on MNIST. On the left we plot the input measurement $\bm{y}$, and the input-adaptive subspace $\mathcal{A}(\bm{y})=\{\bm{x}_{0}(\bm{y}),\bm{w}_{1}(\bm{y}),\bm{w}_{2}(\bm{y})\}$. In the middle, we plot the estimated PPDs obtained from a Gaussian approximation using NPPC, from the KDE approach, and from our PPDE. The projected ground truth posterior sample $\bm{x}$ is marked by a blue cross, and its likelihood under the estimated PPD is shown on the bottom left of the corresponding plot. On the right, we show a few posterior samples out of the 100 used in the KDE approach. These were obtained using an EBM trained on MNIST.
@@
Figure S11: PPD comparisons on CelebA-HQ. The layout is the same as Fig. S10. The samples on the right were obtained using DDNM.
Biological image-to-image translation.
Finally, we test our method on a nonlinear image-to-image translation task. Here, we evaluate PPDE on the dataset from [70], where a microscopic biological specimen imaged with one fluorescent dye is transformed to appear as if it was imaged with another. This “virtual staining” task is of immense importance in biological imaging, as it enables correlative imaging using fewer fluorescent dyes. Figure S9 presents the results of PPDE on this challenging task. The resulting distribution discloses a range of plausible solutions in the projected space corresponding to different numbers of cells. This example demonstrates how PPDE can be used as a quantitative tool for assessing the variance in cell counting applications.
D.3 Quantitative comparisons
Comparison to NPPC.
@@ D.3 Quantitative comparisons
NPPC outputs the posterior mean $\bm{x}_{0}(\bm{y})$, alongside the top $K$ posterior PCs $\bm{W}(\bm{y})$ and variances $\{\sigma_{k}^{2}(\bm{y})\}_{k=1}^{K}$ for a given input image $\bm{y}$. While it was originally proposed as a distribution-free method, the predicted variances can be used to construct a 2D Gaussian approximation of the projected posterior covariance. Let $\bm{\Sigma}(\bm{y})$ denote the diagonal matrix that has $[\sigma_{1}^{2}(\bm{y}),\sigma_{2}^{2}(\bm{y})]^{\top}$ along its main diagonal. Then, the projected posterior distribution can be approximated as
$p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})\approx\mathcal{N}(\bm{v};0,\bm{\Sigma}(\bm{y})).$ (S9)
@@
We compare our results to this baseline in Table S1, showing a significant improvement in sample negative log likelihood (NLL) in favor of PPDE.
@@
Table S1: Negative log likelihood comparison. NPPC and our method were computed on the entire test set. The KDE approach was computed for 100 test images due to computational constraints.
@@
Comparison to posterior sampling + KDE.
The aim of our density estimator is to estimate the PPD in one or two dimensions. Given a method that can sample from the posterior (e.g., DDNM [72]), a possible baseline approach to achieve this goal is as follows: (i) Generate $N$ samples from the posterior $\{\bm{x}_{i}\}$; (ii) Compute PCA using $\{\bm{x}_{i}\}$ and project the samples onto the space spanned by the top two PCs; (iii) Estimate the PPD in 2D from the projected samples $\{\bm{v}_{i}\}$ using standard techniques such as kernel density estimation (KDE). Comparisons against this baseline are reported in Table S1. For producing the KDE samples from the posterior we have used: a conditional EBM for MNIST, DPS [10] [^2] For DPS we have used the public model trained on FFHQ (and not on CelebA). Therefore, the model is not expected to produce accurate samples. The comparison to the PPD it produces is presented here only as an additional rough baseline. for face inpainting [^3] We also experimented with MAT [33], which gave inferior results., and DDNM for face colorization, selecting the best sampler according to their performance and available trained models. For each task/dataset we randomly chose 100 test images and sampled 100 posterior samples per image. Compared to this (impractical) baseline, our technique is orders of magnitude faster as this approach requires costly repeated posterior sampling to estimate the projected density with enough accuracy. Figures S10 and  S11 present visual examples of the approximated PPD with NPPC, KDE, and PPDE. PPDE exhibits superior NLL on average while being extremely faster than KDE, and only slightly slower than NPPC.
E Discussion and conclusion
@@
Acknowledgements
This research was partially supported by the Israel Science Foundation (grant no. 2318/22), by a gift from Elbit Systems, and by the Ollendorff Minerva Center, ECE faculty, Technion.
Appendix A Toy model
A.1 Problem setting
@@
$p_{\mathbf{x}}(\bm{x})=\sum_{\ell=1}^{L}\pi_{\ell}\cdot\varphi(\bm{x};\bm{\mu}_{\ell},\bm{\Sigma}_{\ell}),$ (S10)
@@
$\begin{array}[]{lll}\bm{\mu}_{1}=(-1,1)^{T}&\Sigma_{1}=\begin{pmatrix}0.05&-0.01\\ -0.01&0.025\end{pmatrix}&\pi_{1}=0.15\\ \bm{\mu}_{2}=(1,1)^{T}&\Sigma_{2}=\begin{pmatrix}0.05&0.01\\ 0.01&0.025\end{pmatrix}&\pi_{2}=0.15\\ \bm{\mu}_{3}=(0,0)^{T}&\Sigma_{3}=\begin{pmatrix}0.02&0\\ 0&0.03\end{pmatrix}&\pi_{3}=0.15\\ \bm{\mu}_{4}=(-0.7,-1)^{T}&\Sigma_{4}=\begin{pmatrix}0.15&-0.04\\ -0.04&0.04\end{pmatrix}&\pi_{4}=0.15\\ \bm{\mu}_{5}=(0,-1.2)^{T}&\Sigma_{5}=\begin{pmatrix}0.15&0\\ 0&0.025\end{pmatrix}&\pi_{5}=0.15\\ \bm{\mu}_{6}=(0.7,-1)^{T}&\Sigma_{6}=\begin{pmatrix}0.15&0.04\\ 0.04&0.04\end{pmatrix}&\pi_{6}=0.15.\\ \end{array}$ (S11)
@@
A.2 Analytical posterior distribution
Assuming the setting in Sec. A.1, here we derive the analytical posterior distribution. Specifically, we show that the posterior is also a GMM, and that it has the form
$p_{\mathbf{x}|\mathbf{y}}(\bm{x}|\bm{y})=\sum_{\ell=1}^{L}\tilde{\pi}_{\ell}(\bm{y})\cdot\varphi(\bm{x};\tilde{\bm{\mu}}_{\ell}(\bm{y}),\tilde{\bm{\Sigma}}_{\ell}(\bm{y})),$ (S12)
@@ A.2 Analytical posterior distribution
$\tilde{\pi}_{\ell}(\bm{y})$ $=\frac{\pi_{\ell}\varphi(\bm{y};\bm{\mu}_{\ell},\bm{\Sigma}_{\ell}+\sigma_{n}^{2}\bm{I})}{\sum_{\ell^{\prime}=1}^{L}\pi_{\ell^{\prime}}\varphi(\bm{y};\bm{\mu}_{\ell^{\prime}},\bm{\Sigma}_{\ell^{\prime}}+\sigma_{n}^{2}\bm{I})}.$ (S13)
@@
$p_{\mathbf{x}}(\bm{x})=\sum_{\ell=1}^{L}p_{\mathbf{x}|\textnormal{c}}(\bm{x}|\ell)p_{\textnormal{c}}(\ell).$ (S14)
@@
$p_{\mathbf{x}|\mathbf{y}}(\bm{x}|\bm{y})=\sum_{\ell=1}^{L}p_{\mathbf{x}|\mathbf{y},\textnormal{c}}(\bm{x}|\bm{y},\ell)p_{\textnormal{c}|\mathbf{y}}(\ell|\bm{y}).$ (S15)
The term $p_{\textnormal{c}|\mathbf{y}}(\ell|\bm{y})$ in Eq. (S15) can be written using Bayes’ rule as
@@
$=\frac{\pi_{\ell}\varphi(\bm{y};\bm{\mu}_{\ell},\bm{\Sigma}_{\ell}+\sigma_{n}^{2}\bm{I})}{\sum_{\ell^{\prime}=1}^{L}\pi_{\ell^{\prime}}\varphi(\bm{y};\bm{\mu}_{\ell^{\prime}},\bm{\Sigma}_{\ell^{\prime}}+\sigma_{n}^{2}\bm{I})}.$ (S16)
Moreover, to explicitly express $p_{\mathbf{x}|\mathbf{y},\textnormal{c}}(\bm{x}|\bm{y},\ell)$ in Eq. (S15), we note that $\mathbf{x}$ and $\mathbf{y}$ are jointly Gaussian given c, because under a particular choice $\textnormal{c}=\ell$, the random vector y is the sum of two independent Gaussian vectors, $\mathbf{x}$ and $\mathbf{n}$. Hence, conditioned on the event $\textnormal{c}=\ell$, the joint distribution of $\mathbf{x}$ and $\mathbf{y}$ is given by
$p_{\mathbf{x},\mathbf{y}|\textnormal{c}}(\bm{x},\bm{y}|\ell)$ $=\mathcal{N}\left(\begin{pmatrix}\bm{\mu}_{\ell}\\ \bm{\mu}_{\ell}\end{pmatrix},\begin{pmatrix}\bm{\Sigma}_{\mathbf{x}\mathbf{x}|\textnormal{c}}(\ell)&\bm{\Sigma}_{\mathbf{x}\mathbf{y}|\textnormal{c}}(\ell)\\ \bm{\Sigma}_{\mathbf{x}\mathbf{y}|\textnormal{c}}(\ell)&\bm{\Sigma}_{\mathbf{y}\mathbf{y}|\textnormal{c}}(\ell)\end{pmatrix}\right)$
$=\mathcal{N}\left(\begin{pmatrix}\bm{\mu}_{\ell}\\ \bm{\mu}_{\ell}\end{pmatrix},\begin{pmatrix}\bm{\Sigma}_{\ell}&\bm{\Sigma}_{\ell}\\ \bm{\Sigma}_{\ell}&\bm{\Sigma}_{\ell}+\sigma_{n}^{2}\bm{I}\end{pmatrix}\right).$ (S17)
From the joint distribution in Eq. (S17), the conditional distribution of $\mathbf{x}$ given $\mathbf{y}$ and c is derived as
$p_{\mathbf{x}|\mathbf{y},\textnormal{c}}(\bm{x}|\bm{y},\ell)=\varphi(\bm{x};\tilde{\bm{\mu}}_{\ell}(\bm{y}),\tilde{\bm{\Sigma}}_{\ell}(\bm{y})),$ (S18)
with $\tilde{\bm{\mu}}_{\ell}(\bm{y})$ and $\tilde{\bm{\Sigma}}_{\ell}(\bm{y})$ defined as in Eq. (A.2).
A.3 Projected posterior distribution
Given that the posterior distribution is a GMM as in Eq. (S12), it is straightforward to derive the projected posterior distribution (PPD) over any arbitrary 1D subspace defined by a center $\bm{x}_{0}$ and a normalized direction $\bm{w}$. This is done by projecting each of the Gaussians individually, i.e., the PPD for $v=\bm{w}^{T}(\bm{x}-\bm{x}_{0})$ is given by
$p_{\textnormal{v}|\mathbf{y}}(v|\bm{y})=\sum_{\ell=1}^{L}\tilde{\pi}_{\ell}(\bm{y})\cdot\varphi(v;\bm{w}^{T}(\tilde{\bm{\mu}}_{\ell}(\bm{y})-\bm{x}_{0}),\bm{w}^{T}\tilde{\bm{\Sigma}}_{\ell}(\bm{y})\bm{w}).$ (S19)